OpenStudy (anonymous):

Write the expression -3x squared + 2y squared + 5xy - 2y + 5x squared - 3y squared in simplest form. Then, answer the following questions using complete sentences.

OpenStudy (anonymous):

a. How many terms are in the simplified expression? b. How many of the terms in the simplified expression are negative?

OpenStudy (unimatix):

\[-3x^{2}+2y ^{2}+5xy-2y+5x^{2}-3y ^{2}\] Is this what you have?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 32

OpenStudy (anonymous):

what is the answer to x = 2, y = -1

OpenStudy (anonymous):

and also x = 0, y = 2.5

OpenStudy (texaschic101):

-3x^2 + 2y^2 + 5xy - 2y + 5x^2 - 3y^2 combine like terms -3x^2 + 5x^2 = 2x^2 2y^2 - 3y^2 = -y^2 5xy -- no like terms with this -2y -- no like terms put it all together.. 2x^2 - y^2 + 5xy - 2y -- simplified form you have 4 terms in the simplified form... 2x^2 is a term, -y^2 is also a term....5xy is a term...and -2y is a term you have 2 negative terms.....-y^2 and -2y

OpenStudy (anonymous):

let me simplify it match up column A with column B: x = 2, y = -1 14 2. x = 0, y = 2.5 1.665 3. x = -1, y = -3 -11.25 4. x = 0.5,y=negative 1/10 0.44 5. x = 3/4, y =2/5 -1 6. x = √2, y = √2 9.17

OpenStudy (texaschic101):

I don't understand what you are supposed to do with these...your just given x and y points....where is the rest of this problem ?

OpenStudy (anonymous):

that is what i was given and you match them up

OpenStudy (anonymous):

the best you can

OpenStudy (texaschic101):

are you supposed to sub in your x and y points into that first question ?

OpenStudy (anonymous):

no they are two different questions

OpenStudy (texaschic101):

then the matching thing makes no sense ....I am sorry but I don't understand.

OpenStudy (anonymous):

for number 4 the negative 1/10 is a negative fraction

OpenStudy (anonymous):

neither do i that is why i put this question on here

OpenStudy (texaschic101):

Oh crap...I lost track of time....I have to be somewhere in 20 minutes and I am not even ready. You are going to have to tag someone...maybe they will have better luck. Again, I am sorry

OpenStudy (anonymous):

please help me out on my second question

OpenStudy (unimatix):

Is the second question a sub question of the first one?

OpenStudy (texaschic101):

thats is what I thought...but he said it was a separate question....not related

OpenStudy (anonymous):

yea it was i solved the question sorry lol

OpenStudy (texaschic101):

good to hear you solved it :) way to go